You're trying to determine whether to expand your business by building a new manufacturing plant. The plant has an installation cost of $12 million, which will be depreciated straight-line to zero over its four-year life. If the plant has projected net income of $1,856,100, $1,909,400, $1,877,800, and $1,331,300 over these four years, what is the project's average accounting return (AAR)? (Do not round your intermediate calculations. Round the final answer to 2 decimal places. Omit % sign in your response.) 
Answer #1

Initial investment = $12,000,000

Useful life = 4 year

Average investment = Initial investment x 1/12

= 12,000,000 x 1/2

= $6,000,000

Total net income for 4 year = 1,856,100 + 1,909,400 + 1,877,800 + 1,331,300

= $6,974,600

Average annual Net income = Total net income for 4 year/4

= 6,974,600/4

= $1,743,650

Average accounting return = Average annual Net income/Average investment

= 1,743,650/6,000,000

= 29.06